Perform the indicated operation.
g(n)= n^2+4+2n
h(n)= -3n+2
find (g * h)(1)
thanks.

Answer by rapaljer(4671)

g(1)= 1+4+2=7

h(1) = -3+2=-1

(g*h)(1) = g(1)*h(1) = 7*(-1)=-7

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus